SEPARATED NETS IN EUCLIDEAN SPACE AND JACOBIANS OF BILIPSCHITZ MAPS

DIMITRI BURAGO BRUCE KLEINER

Abstract. We show that there are separated nets in the Euclidean plane which are not biLipschitz equivalent to the integer lattice. The argument is based on the construction of a continuous function which is not the Jacobian of a biLipschitz map.